# Index Terms: smartphone, vibration, tactile, perceptionI Introduction

### 14 Jan 2024 v1

@@ II-C Statistical Analyses
Differences in perception data at the index finger and big toe (H1 and F from Fig. 2) are calculated using paired single-sided t-tests with Bonferroni correction for the smartphone and tuning fork and a single-sided Wilcoxon signed rank test (which is the nonparametric equivalent) for the monofilament.
Differences in perception data between younger adults and older adults at the index finger are calculated using unpaired single-sided t-tests with Bonferroni correction for the smartphone and tuning fork and a single-sided Wilcoxon rank sum test with Bonferroni correction (which is the nonparametric equivalent) for the monofilament. Differences in perception data between younger adults and older adults at big toe are calculated in the same manner.
@@ III-A Discrimination Between Hands and Feet
To investigate the resolution limits of our smartphone-based tool, we first sought to determine whether our tool could detect known trends in vibration perception differences of different body locations in healthy humans. As shown in Fig. 3A, our participants had smartphone threshold value of “hapticIntensity” () at the index finger and “hapticIntensity” at the big toe. From our statistical analysis, we found that our smartphone-based tool detected a statistically significant VPT difference between the index finger and big toe (). We also found that our smartphone-based tool detected a large effect size between the index finger and the big toe ().
@@ III-B Discrimination Between Younger and Older Adults
We also investigated the smartphone’s ability to measure known effects of age on vibrotactile perception (Fig. 4A). From our statistical analyses, we found that the smartphone could not discriminate between younger ( “hapticIntensity” ) and older ( “hapticIntensity”) adults at the index finger location () but could discriminate between younger ( “hapticIntensity”) and older ( “hapticIntensity” ) adults at the big toe (). We also found that the smartphone detected a small effect size between the index fingers of the younger and older adult (). There was a large effect size between the big toes of the younger and older adult ().
We ran the same statistical tests on the tuning fork exam data for clinical comparison. As shown in Fig. 4B, we found that the tuning fork exam could not discriminate between younger ( s) and older ( s ) adults at the index finger location (). However, the tuning fork exam could discriminate between younger ( s) and older ( s) adults at the big toe location (). We also found that the tuning fork detected a moderate effect size between the index fingers of the younger and older adult (). There was a large effect size between the big toes of the younger and older adult ().
@@ III-C Correlation Between Smartphone VPT and Clinical Measurements
The correlations between the various touch sensitivity tools are all positive as expected, with the strongest correlation existing between the smartphone and the tuning fork. Both the smartphone and tuning fork measure vibrotactile ability, so we expect those to be more closely correlated than the monofilaments (which measure force, not vibration perception). However, the correlations are quite weak. One possible reason for this result is that the tuning fork and monofilament perception values are subject to inconsistencies, such as variation in how hard the experimenter struck the tuning fork and how hard the experimenter pressed the monofilaments. Another possible reason is that we chose locations that did not yield much spread in the VPT data. To better assess this idea, we calculated correlations for the smartphone VPT vs. tuning fork perception time when one focuses solely on the index finger and foot locations. The index finger is a convenient and common vibration perception testing location and the foot is important testing location for clinical applications. We noticed that when one isolates the index finger and foot location, there’s a much stronger correlation between the phone and the clinical tools. As shown in Fig. 5D, we found a statistically significant, and moderate, negative correlation between the smartphone-based VPT and the tuning fork VPT (, ). As shown in Fig. 5E, we found a statistically significant, and moderate, positive correlation between the smartphone-based VPT and the monofilament threshold (, ). As shown in Fig. 5F, we found a statistically significant, and moderate, negative correlation between the monofilament threshold and the tuning fork perception time (, ).
We also visualized how vibration perception changes across different body parts with the phone and the tuning fork in Fig. 6. As one would expect, the trend across body parts for for the smartphone is inverted from the trend across body parts for the tuning fork. The one exception is the W2 location, and we believe this may be due to a confounding factor of the experimental setup. Later in the the data collection, we realized that many participants could feel W2 vibrations in their fingers as well as through the wrist because of how vibrations from the phone are propagated through the pillow to the fingers in the W2 configuration (Fig. 2). This would have caused the W2 smartphone-based VPTs to be more similar to the hand smartphone-based VPTs.
@@ Acknowledgments
The authors thank Bryce N. Huerta and Alexis J. Lowber for their contributions to the early iPhone application development and Adeyinka E. Adenekan, Ph.D. for insightful discussions regarding the data visualization.

### 25 Jan 2024 v2

@@ II-C Statistical Analyses
Differences in thresholds measured at the index finger and big toe (H1 and F from Fig. 2) are analyzed using paired single-sided t-tests with Bonferroni correction for the smartphone and tuning fork and a single-sided Wilcoxon signed rank test (the nonparametric equivalent) for the monofilament.
Differences in thresholds measured between younger adults and older adults at the index finger are analyzed using unpaired single-sided t-tests with Bonferroni correction for the smartphone and tuning fork and a single-sided Wilcoxon rank sum test with Bonferroni correction (the nonparametric equivalent) for the monofilament. Differences in thresholds measured between younger adults and older adults at big toe are analyzed in the same manner.
@@ III-A Discrimination Between Hands and Feet
To investigate the resolution limits of our smartphone-based tool, we first sought to determine whether our tool could detect known trends in vibration perception of different body locations in healthy humans. As shown in Fig. 3A, our participants had a smartphone threshold value of hapticIntensity () at the index finger and hapticIntensity at the big toe. From our statistical analysis, we found that our smartphone-based tool detected a statistically significant VPT difference between the index finger and big toe (). We also found that our smartphone-based tool detected a large effect size between the index finger and the big toe ().
@@ III-B Discrimination Between Younger and Older Adults
We also investigated the smartphone’s ability to measure known effects of age on vibrotactile perception (Fig. 4A). From our statistical analyses, we found that the smartphone could not discriminate between younger ( hapticIntensity) and older ( hapticIntensity) adults at the index finger location () but could discriminate between younger ( hapticIntensity) and older ( hapticIntensity) adults at the big toe (). We also found that the smartphone detected a small effect size between the index fingers of the younger and older adults (). There was a large effect size between the big toes of the younger and older adults ().
We ran the same statistical tests on the tuning fork exam data for clinical comparison. As shown in Fig. 4B, we found that the tuning fork exam could not discriminate between younger ( s) and older ( s) adults at the index finger location (). However, the tuning fork exam could discriminate between younger ( s) and older ( s) adults at the big toe location (). We also found that the tuning fork detected a moderate effect size between the index fingers of the younger and older adults (). There was a large effect size between the big toes of the younger and older adults ().
@@ III-C Correlation Between Smartphone VPT and Clinical Measurements
The correlation between the smartphone and tuning fork is negative as expected. The correlation between the smartphone and monofilaments is positive as expected. Both the smartphone and tuning fork measure vibrotactile ability, so we expect those to be more closely correlated than the monofilaments (which measure force, not vibration perception). However, the correlations are quite weak. One possible reason for this result is that the tuning fork and monofilament perception values are subject to inconsistencies, such as variation in how hard the experimenter struck the tuning fork and how hard the experimenter pressed the monofilaments. Another possible reason is that we chose locations that did not yield much spread in the VPT data. To better assess this idea, we calculated correlations for the smartphone VPT vs. tuning fork perception time when one focuses solely on the index finger and foot locations. The index finger is a convenient and common vibration perception testing location and the foot is an important testing location for clinical applications. We noticed that when one isolates the index finger and foot location, there is a much stronger correlation between the phone and the clinical tools. As shown in Fig. 5D, we found a statistically significant, and moderate, negative correlation between the smartphone-based VPT and the tuning fork VPT (, ). As shown in Fig. 5E, we found a statistically significant, and moderate, positive correlation between the smartphone-based VPT and the monofilament threshold (, ). As shown in Fig. 5F, we found a statistically significant, and moderate, negative correlation between the monofilament threshold and the tuning fork perception time (, ).
We also visualized how vibration perception changes across different body parts with the phone and the tuning fork in Fig. 6. As one would expect, the trend across body parts for the smartphone is inverted from the trend across body parts for the tuning fork. The one exception is the W2 location, and we believe this may be due to a confounding factor of the experimental setup. Later in the data collection, we realized that many participants could feel W2 vibrations in their fingers as well as through the wrist because of how vibrations from the phone are propagated through the pillow to the fingers in the W2 configuration (Fig. 2). This would have caused the W2 smartphone-based VPTs to be more similar to the hand smartphone-based VPTs.
@@ Acknowledgments
The authors thank the participants for their time and efforts in providing us with a rich dataset, J. Mucini, M.S., OTR/L for her clinical advice regarding the monofilament and tuning fork protocols, Bryce N. Huerta, M.S. and Alexis J. Lowber, M.S. for their contributions to the early iPhone application development, and Adeyinka E. Adenekan, Ph.D. for insightful discussions regarding the data visualization.